Madsen Motors's bonds have 5 years remaining to maturity. Interest is paid annually, they have a $1,000 par value, the coupon interest rate is 7%, and the yield to maturity is 8%. What is the bond's current market price? Round your answer to the nearest cent.
Answer #1

Annual coupon=1000*7%=70

Hence current price=Annual coupon*Present value of annuity factor(8%,5)+$1000*Present value of discounting factor(8%,5)

=70*3.99271004+1000*0.680583197

=$960.07(Approx)

NOTE:

1.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=70[1-(1.08)^-5]/0.08

=70*3.99271004

2.Present value of discounting factor=1000/1.08^5

=1000*0.680583197